# Beyond Linearity and Time-homogeneity: Relational Hyper Event Models with Time-Varying Non-Linear Effects

### v2

@@ 3 Simulation study
Depending to the simulation setup, covariates effects are selected as LE, TVE, NLE, and TVNLE. This study was designed with four objectives in mind. First, we aim to demonstrate that when the true underlying effect is linear, both NLE and TVNLE specifications are flexible enough to recover linearity as a special case, as discussed in Section 2.2. Second, we highlight that when the true effect varies over time, fitting a linear model fails to capture the temporal dynamics accurately. Third, we show the potential of the proposed joint TVNLE estimation procedure in capturing complex covariate effects. Finally, we emphasize that using only time-varying or only non-linear models may be insufficient when the data-generating process involves both temporal and non-linear components.
@@
Linear models are recovered by penalized non-linear models.
We generate synthetic hyperevent data according to the following underlying model:
$\lambda(t,I)=-\log\left(\sqrt{\text{subrep}^{1}(t,I)}\right)+\log\left(\sqrt{\text{subrep}^{2}(t,I)}\right)+\overline{x}(I)^{2}-0.5\cdot|I|.$ (10)
Simulated hyperevents represent undirected meetings involving multiple interacting actors, without distinguishing between senders and receivers. Consequently, the intensity function depends solely on the set $I$ of participants. The terms $\text{subrep}^{1}(t,I)$ and $\text{subrep}^{2}(t,I)$ are endogenous covariates capturing first- and second-order subset repetition, respectively, as defined in Equation (4). The variable $\overline{x}(I)$ is an exogenous covariate computed as the average of a Gaussian-distributed individual feature across all members of $I$. Finally, $|I|$ denotes the size of the meeting, included in the model with a negative contribution to penalize larger events. If $\overline{x}(I)^{2}$ is used as the covariate — instead of $\overline{x}(I)$ — its effect remains linear within the model specification, as the transformation is applied to the covariate before modeling.
Figure 1 aggregates both linear and non-linear estimates from multiple replications using inverse-variance weighting (Hartung et al., 2011).
@@
Linear models fail to capture time-varying effects.
We simulate hyperevent data from the true underlying model with time-varying effects,
$\lambda(t,I)=\underbrace{-10\cdot t}_{\beta_{1}(t)}\cdot\log\left(\sqrt{\text{subrep}^{1}(t,I)}\right)+\underbrace{10\cdot t}_{\beta_{2}(t)}\cdot\log\left(\sqrt{\text{subrep}^{2}(t,I)}\right)+\underbrace{10\cdot t}_{\beta_{3}(t)}\cdot\overline{x}(I)^{2}\underbrace{-0.1}_{\beta_{4}}\cdot|I|.$ (11)
Here, $\beta_{j}(t)$ for $j=1,2,3$ are time-varying functions. By contrast, the true effect of event size $\beta_{4}<0$ is time-invariant and linear. Figure 2 shows that the linear model fails to capture the time-varying effect of this variable. By contrast, the TVE model successfully identifies it. The TVNLE model also correctly identifies the time-varying — but, for each time point, linear — effect. As it can be seen, for time equal to zero, the effect does not increase with the covariate (constant color on the vertical line at $t=0.0$), while the vertical color gradient becomes steeper as time increases.
The potential of the joint time-varying non-linear estimation.
In the previous analysis, we assumed knowledge of the true non-linear transformation of the covariate, applying it explicitly in the model. This is clearly unrealistic in practice. In this third section, we therefore remove the known square transformation of $\overline{x}(I)$ — which we know to be the true form — and instead allow the transformation to be learned from the data itself, permitting it to vary over time.
Figure 3 shows the true centered effect (Top Left), i.e. $f^{\text{TVNLE}}\left(\overline{x}(I),t\right)=+10\cdot t\cdot\overline{x}(I)^{2}$ and the estimated effect $\hat{f}^{\text{TVNLE}}\left(\overline{x}(I),t\right)$ (Top Right). The color shading represents the effect, plotted with time $t$ on the x-axis and the values of the hyperevent covariate $\overline{x}(I)$ on the y-axis. We observe that the estimates successfully capture both the variation over time and the variation across covariate values. The image was generated by predicting the log-hazard contributions for 30 pairs of time and covariate values, uniformly sampled from the range of the synthetic data.
@@
Figure 3: The potential of joint time-varying non-linear estimation and the failure of time-varying and non-linear models. The four plots display results from simulated data based on the model in Equation 11. The two upper plots in Figure 3 show centered true effect and the estimated effect, respectively. The color shading represents the effect $f^{\text{TVNLE}}(\overline{x}(I),t)$, with time $t$ on the x-axis and the values of the hypercovariate $\overline{x}(I)$ on the y-axis. The two lower plots in Figure 3 show interpolated values of the estimated contributions to the log-rate function, obtained by fitting a model with only time-varying effects and non-linear effects, respectively. The contribution to the log-rate with TVE consists of the TVE itself multiplied by the value of the covariate at the corresponding time.
Time-varying and non-linear models fail to capture simultaneous time-varying non-linear effects.
Figure 3 Bottom Left and Bottom Right displays interpolated values of the estimated contributions to the log-rate function, obtained by fitting a model with time-varying effects and non-linear effects only. It is important to note that when considering TVE, the log-hazard contribution corresponds to the time-varying effect multiplied by the value of the covariate at the corresponding time. We observe that the TVE contribution incorrectly identifies a decrease in the value of the contribution for low covariate values towards the end of the time window. In contrast, the NLE model fails to capture the temporal variation in the data.
@@ 4 Empirical Application to Coauthorship Citation Networks
We applied our RHEM extension to the DBLP-Citation-network V14 dataset, restricted to journal articles, derived from the Aminer Citation Network (Tang et al., 2008). In this context, a relational hyperevent is defined as the formal publication of a scientific work, marking the moment when a group of authors presents their research alongside a set of citations listed in the reference section. Our analysis is specifically focused on journal publications, constructing an event list that includes only these types of works. Furthermore, among the cited works, we also restrict our selection to journal publications to maintain consistency within the dataset. The resulting event sequence comprises $n=1,416,353$ publication events spanning from 1939 to 2023.
@@
Here, $j_{m}$ represents the journal article published at time $t_{m}$ by a group of authors in $I_{m}\subseteq V^{I}_{t_{m}}$ citing journal articles in $J_{m}\subseteq V^{J}_{t_{m}}$. This dataset was previously analyzed using a standard linear RHEM approach (Lerner et al., 2025). We selected it as an empirical test case to demonstrate the advantages of our proposed model over the current state-of-the-art methodology as represented by the cited study. Code to preprocess and filter the data is publicly available. [^3] https://github.com/juergenlerner/eventnet/tree/master/data/scientific_networks/aminer_2023 The counting process $N(t,I,J)$ tracks the cumulative number of papers authored by group $I$ that cite works in set $J$ up to time $t$. This approach examines how endogenous relational mechanisms — particularly prior scientific collaborations or citation patterns among authors — influence the publication rate of journal articles.
@@
The inclusion of time-varying effects is particularly pertinent in this analysis, given that publication events are not uniformly distributed: over 87% of the publications occurred after 2000. By including in the model formulation an explicit reference to the time of the event, interacting with the covariate evaluated at the time of interest, our technique allows to account for this temporal imbalance. Moreover, this empirical application carries significant social implications. Scientific networks provide insights into the factors motivating different authors to collaborate, revealing the complex dynamics that drive co-authorship and knowledge exchange. This presents a valuable opportunity to deepen our understanding of the evolving dynamics within complex social networks, particularly within the scientific domain. These networks consist of interconnected authors and papers, but they can also include inventors, awardees, investors, publications, grants, and patents, along with their intricate relationships. Such networks are continually expanding, embodying social phenomena like the exchange of funding, knowledge, and reputation. There is no inherent justification for assuming that this influence is linear or homogeneous over time, as collaboration patterns evolve in response to shifts in research interests, available resources, and community structures, as demonstrated in various approaches to analyzing co-authorship networks (Hoekman et al., 2010; Kwiek, 2021).
@@ 4.1 Drivers of Scientific Collaboration and Impact
Using the case-control partial likelihood inference framework with GAMs, as explained in Section 2.3, we sampled one possible but unobserved event from the risk set for each publication event. For each group of authors $I$ and the corresponding group of cited papers $J$ at time $t$, we selected a pair $(I^{\ast},J^{\ast})\in\binom{V^{I}_{t}}{|I|}\times\binom{V^{J}_{t}}{|J|}$, ensuring both groups have the same size in the hyper-event and the non-hyperevent. Hyperedge covariates were generated using the open-source software eventnet Lerner and Lomi (2020, 2023), [^4] https://github.com/juergenlerner/eventnet which allows sampling a specified number of non-event hyperedges related to each observed publication event, and calculating covariates for both observed events and sampled non-events. In our analysis, we sample one non-event ($m=1$).
In this empirical application, relational hyperevents represent citations. Therefore, we replace the generic term “action” in Equation (3) with “citation”. We focus on a subset of the hyperedge covariates studied in Lerner et al. (2025) and, following that work, incorporate a weight factor $\omega$ to model the temporal decay of past events. Most of the statistics used can be defined by adapting and combining the function “citation”. Specifically, we introduce the following adaptations, based on the superscript referring to either authors or papers (aut or pap):
@@
    Paper-paper citations:
@@
    Author-Author citations:
@@
where $\omega(t-t_{m})=\exp\{-\left(t-t_{m}\right)\frac{\log 2}{T_{\frac{1}{2}}}\}$ and $T_{\frac{1}{2}}>0$ is the half-life period (Lerner et al., 2013).
As we show that “action” defined in Equation (3) serves as a basic block for subset-repetition “subrep” defined in Equation (4), all the previously listed adaptations play a similar role. Specifically, we focus our attention on the following hyperedge covariates involving authors’ network, papers’ network and their interconnection (see Lerner et al. (2025) for a more detailed description, graphical illustration, and numerical examples).
@@
To improve model fitting stability, we modified the network statistics by using a different approach from the square-root transformation proposed in Lerner et al. (2025). Similar to their goal, our aim is to reduce the effect of outliers on the model. To achieve this, each covariate—whether for observed events or non-events—is transformed using the following formula:
$\dot{x}^{k}=1-\exp\left(-\frac{x^{k}}{c^{k}}\right)$ (12)
Here, $x^{k}$ is a value of the $k$-th covariate, observed for either events or non-events. The parameter $c^{k}$ is selected through an optimization procedure. This procedure finds the value of $c^{k}$ that brings the transformed covariate as close as possible to a uniform distribution, based on minimizing the Kolmogorov–Smirnov statistic. Additionally, we applied a transformation to the timing of the events. Since time is evaluated in the same way for both events and non-events, we used the empirical cumulative distribution function (ECDF) of the event times to transform them into a uniform distribution:
@@
where $\text{ecdf}_{t}$ is the ECDF of the event times. Although these transformations may seem to make interpretation harder, after fitting we map values back to the original scale and interpret the results in terms of the original covariates.

### v3

@@ 4 Simulation study
Depending to the simulation setup, covariates effects are selected as LE, TVE, NLE, and TVNLE. This study was designed with four objectives in mind. First, we aim to demonstrate that when the true underlying effect is linear, both NLE and TVNLE specifications are flexible enough to recover linearity as a special case – as discussed in Section 3.2. Second, we highlight that when the true effect varies over time, fitting a linear model fails to capture the temporal dynamics accurately. Third, we show the potential of the proposed joint TVNLE estimation procedure in capturing complex covariate effects. Finally, we emphasize that using only time-varying or only non-linear models may be insufficient when the data-generating process involves both temporal and non-linear components.
@@
4.1 Linear models are recovered by penalized non-linear models.
In this section, we simulate data from a rate that is linear in one of the covariates, $\bar{x}(I)^{2}$, but we fit it using a non-linear effect, $f(\bar{x}(I)^{2})$. The aim is to assess whether the non-linear model overfits the effect. More specifically, the underlying true data-generating process has the following intensity function:
$\log\lambda(t,I)=-\log\left(\sqrt{\text{subrep}^{1}(t,I)}\right)+\log\left(\sqrt{\text{subrep}^{2}(t,I)}\right)+\overline{x}(I)^{2}-0.5\cdot|I|.$ (9)
Simulated events are undirected hyperevents, i.e., meetings involving multiple actors simultaneously, without distinguishing between senders and receivers.. Consequently, the intensity function depends solely on the set $I$ of participants. The terms $\text{subrep}^{1}(t,I)$ and $\text{subrep}^{2}(t,I)$ are endogenous covariates capturing first- and second-order subset repetition, respectively, as defined in Equation (4). The variable $\overline{x}(I)$ is an exogenous covariate computed as the average of the individual feature across all members of $I$. Finally, $|I|$ denotes the size of the meeting, included in the model with a negative contribution to penalize larger events. If $\overline{x}(I)^{2}$ is used as the input covariate – instead of $\overline{x}(I)$ – its effect remains linear within the model specification, as the transformation is applied to the covariate before modeling. Note that the coefficients for first- and second-order subset repetition are chosen with opposite signs. If both were positive, the simulation would become unstable due to the nature of this covariate.
Figure 5 aggregates both linear and non-linear estimates from multiple replications using inverse-variance weighting (hartung2011statistical).
@@
4.2 Linear models fail to capture time-varying effects.
In this section we simulate data with various time-varying effects $\gamma_{j}(t)$. The aim is to show that the traditional REM model fails to capture this aspect, whereas our REM extension correctly captures its time-varying nature. We simulate hyperevent data from the true underlying model with time-varying effects,
$\log\lambda(t,I)=\underbrace{-10\cdot t}_{{\gamma_{1}(t)}}\cdot\log\left(\sqrt{\text{subrep}^{1}(t,I)}\right)\underbrace{+10\cdot t}_{{\gamma_{2}(t)}}\cdot\log\left(\sqrt{\text{subrep}^{2}(t,I)}\right)+\underbrace{+10\cdot t}_{{\gamma_{3}(t)}}\cdot\overline{x}(I)^{2}-0.1\cdot|I|.$ (10)
Here, $\gamma_{k}(t)$ for $k=1,2,3$ are time-varying functions. By contrast, the true effect of event size is time-invariant and linear. More specifically, we chose a time-varying function that is linear in time, i.e. $\gamma_{1}(t)=-10t,\gamma_{2}(t)=10t,\gamma_{3}(t)=10t$. In principle, it is possible to simulate more complex functions, such as the non-monotonic example in Figure 2. However, the goal here is to show that even with a simple time-varying function, a linear effect fails to capture the true trend. As in Equation 9, the coefficients $\gamma_{1}(t)=-10t$ and $\gamma_{2}(t)=10t$ were chosen to balance each other, preventing the rate from exploding as events accumulate.
Figure 6 shows that the linear model fails to capture the time-varying effect of this variable. By contrast, the TVE model successfully identifies it. The TVNLE model also correctly identifies the time-varying – but, for each time point, linear – effect. As can be seen, for time equal to zero, the effect does not increase with the covariate (constant color on the vertical line at $t=0.0$), while the vertical color gradient becomes steeper as time increases.
4.3 The potential of the joint time-varying non-linear estimation.
In the previous analysis, we assumed knowledge of the true non-linear transformation of the covariate, applying it explicitly in the model. This is clearly unrealistic in practice. In this third section, we therefore do not assume to know the true square transformation of $\bar{x}(I)$ nor its time-varying nature and instead allow the transformation $f$ to be learned from the data itself, permitting it to vary over both $\bar{x}(t)$ and time $t$, i.e., $f(\bar{x}(t),t)$.
Figure 7 a) shows the true centered effect, i.e. $f^{\text{TVNLE}}\left(\overline{x}(I),t\right)=+10\ t\ \overline{x}(I)^{2}$ and b) the estimated effect $\hat{f}^{\text{TVNLE}}\left(\overline{x}(I),t\right)$. The color shading represents the effect, plotted with time $t$ on the x-axis and the values of the hyperevent covariate $\overline{x}(I)$ on the y-axis. We observe that the estimates successfully capture both the variation over time and the variation across covariate values. The image was generated by predicting the log-hazard contributions for 30 pairs of time and covariate values, uniformly sampled from the range of the synthetic data.
a) b)
@@
Figure 7: The potential of joint time-varying non-linear estimation and the failure of time-varying and non-linear models. The four plots display results from simulated data based on the model in Equation 10. The two upper plots, a) and b), show centered true effect and the estimated effect, respectively. The color shading represents the effect $f^{\text{TVNLE}}(\overline{x}(I),t)$, with time $t$ on the x-axis and the values of the hypercovariate $\overline{x}(I)$ on the y-axis. The two lower plots, c) and d) show interpolated values of the estimated contributions to the log-rate function, obtained by fitting a model with only time-varying effects and non-linear effects, respectively. The contribution to the log-rate with TVE consists of the TVE itself multiplied by the value of the covariate at the corresponding time.
4.4 Time-varying and non-linear models fail to capture simultaneous time-varying non-linear effects.
Using the same simulation following the model in (10), this section aims to show that joint time-varying and non-linear estimates are strictly necessary to uncover the underlying structure of the dynamics. Only time-varying or only non-linear effects fail dramatically, leading to wildly different conclusions.
Figure 7 c) and d) display interpolated values of the estimated contributions to the log-rate function, obtained by fitting a model with time-varying effects and non-linear effects only. It is important to note that when considering TVE, the log-hazard contribution corresponds to the time-varying effect multiplied by the value of the covariate at the corresponding time. We observe that the TVE contribution incorrectly identifies a decrease in the value of the contribution for low covariate values towards the end of the time window. In contrast, the NLE model fails to capture the temporal variation in the data.
@@ 5 Empirical Application to Coauthorship Citation Networks
1. Compute covariates for events and non-events using eventnet Using the eventnet software, 1. Extract event information from the raw data. 2. Specify the components required by eventnet to define the risk set. 3. Sample a non-event observation from the risk set. 4. Compute the covariates for both events and non-events. A tutorial and example code for creating an eventnet configuration for publication event data is provided at the following Link. 2. Create the case-control dataset Merge the output of eventnet by event and non-event identifiers so that each row contains information for both the event and the non-event. 3. Transform covariates and event times 1. Apply a monotonic transformation to each covariate $x$: $\dot{x}=1-\exp\left(-\frac{x}{c}\right)$ Notes: Choose $c$ so that the transformed covariate (evaluated for events) is as close as possible to a uniform distribution. Apply the same transformation to both events and non-events. 2. Transform event times $t$ using the empirical cumulative distribution function: $\dot{t}=\mathrm{ecdf}_{t}(t)$ 4. Fit the model using mgcv::gam()
n = nrow(data)
X = cbind(data\$transf_x_event,
               data\$transf_x_non_event)
T = cbind(data\$transf_time,
               data\$transf_time)
unit = rep(1, n)
I = cbind(unit, -unit)
gam(y ~ -1 + te(T, X, by = I),
      family = "binomial")
5. Inference and diagnostics. Interpret the obtained time-varying non-linear contributions. Example: AAS. This effect exhibits a non-monotonic pattern both over time and across covariate values. A detailed interpretation is provided in Figure 3.Refer to caption
Figure 8: Flow chart summarizing the steps required to replicate the analysis.
We applied our RHEM extension to the DBLP-Citation-network V14 dataset, restricted to journal articles, derived from the Aminer Citation Network (tang2008arnetminer). In this context, a relational hyperevent is defined as the formal publication of a scientific work, marking the moment when a group of authors presents their research alongside a set of citations listed in the reference section. Our analysis is specifically focused on journal publications, constructing an event list that includes only these types of works. Furthermore, among the cited works, we also restrict our selection to journal publications to maintain consistency within the dataset. Code to pre-process and filter the data is publicly available. [^3] https://github.com/juergenlerner/eventnet/tree/master/data/scientific_networks/aminer_2023
The resulting event sequence comprises $n=1,416,353$ publication events spanning from 1939 to 2023.
@@
Here, $j_{m}$ represents the journal article published at time $t_{m}$ by a group of authors in $I_{m}\subseteq V^{I}_{t_{m}}$ citing journal articles in $J_{m}\subseteq V^{J}_{t_{m}}$. The events in the empirical data can be represented exactly as in Figure 1, namely as time-stamped directed hyperedges of a two-mode network. There are two types of nodes: authors and cited papers. When a new hyperedge occurs at time $t_{m}$, a new publication $j_{m}$ – authored by a set of authors in $I_{m}$ and citing a set of papers in $J_{m}$ – enters the system and can be cited by other papers at later time points. This two-mode network representation also makes it possible to examine more complex relationships, such as author-cites-paper, paper-cites-paper, and author-cites-author. These three relations are part of a broader set used to compute the covariates defined in Section 5.1. For this reason, labeling this structure simply as a bipartite network in which authors are connected to the papers they cite may be an oversimplification. Specifically, the counting process $N(t,\cdot,\cdot)$ tracks the cumulative number of papers produced by authors citing existing works up to time $t$. This approach allows us to examine how endogenous relational mechanisms – particularly prior scientific collaborations or citation patterns among authors – influence the publication rate of journal articles $\lambda(t,\cdot,\cdot)$. The co-evolution of co-authoring and citation networks was first proposed by (lerner2024relational). In this paper, we relax their assumption of linear and time-homogeneous effects, showing that such flexibility is warranted for these data.
The empirical data under consideration spans the period from 1939 to 2023. Event times are recorded with yearly granularity, so many papers appear as occurring simultaneously. From a theoretical point of view, this does not imply that these events actually occurred at the same moment (indeed, it is likely not the case). However, because we lack information to determine their order, some adjustments are required. First, when building the case-control data set, the sampled non-hyperevent must differ not only from the corresponding hyperevent but also from all hyperevents occurring in the same year. Second, when computing endogenous covariates, publication events from the same year cannot be treated as previously occurred events. In practice, the eventnet software accounts for this by allowing the user to specify the variable that identifies simultaneous events, such as the yearly event time recorded in this data set.
@@
Event times are also not uniformly distributed, which further motivates the inclusion of time-varying effects. Indeed, over 87% of the publications occurred starting from 2000. By including in the model formulation an explicit reference to the time of the event, interacting with the covariate evaluated at the time of interest, our technique allows us to account for this temporal imbalance. Moreover, this empirical application carries significant social implications. Scientific networks provide insights into the factors motivating different authors to collaborate, revealing the complex dynamics that drive co-authorship and knowledge exchange. This presents a valuable opportunity to deepen our understanding of the evolving dynamics within complex social networks, particularly within the scientific domain. These networks consist of interconnected authors and papers, but they can also include inventors, awardees, investors, publications, grants, and patents, along with their intricate relationships. Such networks are continually expanding, embodying social phenomena like the exchange of funding, knowledge, and reputation. There is no inherent justification for assuming that this influence is linear or homogeneous over time, as collaboration patterns evolve in response to shifts in research interests, available resources, and community structures, as demonstrated in various approaches to analyzing co-authorship networks (hoekman2010research; kwiek2021large). Other motivations for this study can be found in Section 2. A flow chart summarizing the steps required to replicate the analysis is provided in Figure 8.
@@ 5.1 Drivers of Scientific Collaboration and Impact
Using the case-control partial likelihood inference framework with GAMs, as explained in Section 3.3, we sampled one possible but unobserved event from the risk set for each publication event. For each group of authors $I$ and the corresponding group of cited papers $J$ at time $t$, we selected a pair $(I^{\ast},J^{\ast})\in\binom{V^{I}_{t}}{|I|}\times\binom{V^{J}_{t}}{|J|}$, ensuring both groups have the same size in the hyper-event and the non-hyperevent. Hyperedge covariates were generated using the open-source software eventnet lerner2020reliability; lerner2023relational, [^4] https://github.com/juergenlerner/eventnet which allows sampling a specified number of non-event hyperedges related to each observed publication event, and calculating covariates for both observed events and sampled non-events. In our analysis, we sample one non-event ($m=1$).
In this empirical application, relational hyperevents represent citations. Therefore, we replace the generic term “action” in Equation (3) with “citation”. We focus on a subset of the hyperedge covariates studied in lerner2024relational and, following that work, incorporate a weight factor $\omega$ to model the temporal decay of past events. Most of the statistics used can be defined by adapting and combining the function “citation”. Specifically, we introduce the following adaptations, based on the superscript referring to either authors or papers (aut or pap):
@@
    Paper-paper citations::
@@
    Author-Author citations::
@@
where $\omega(t-t_{m})=\exp\Big\{-(t-t_{m})\frac{\log 2}{T_{\frac{1}{2}}}\Big\}$ and $T_{\frac{1}{2}}>0$ is the half-life period (lerner2013modeling), set to three years in this empirical application, following lerner2024relational.
As we show that “action” defined in Equation (3) serves as a basic block for subset-repetition “subrep” defined in Equation (4), all the previously listed adaptations play a similar role. Specifically, we focus our attention on the following hyperedge covariates involving authors’ network, papers’ network and their interconnection (see lerner2024relational for a more detailed description, graphical illustration, and numerical examples).
@@
An implicit assumption in fitting any regression model using maximum likelihood estimation is that the covariate processes have finite second moments. Practically speaking, this means that for improved model fitting stability, the covariates should not have major outliers. lerner2024relational employ a square-root transformation of the non-negative covariates. As this only provides partial stabilization, we modified the network statistics using a different approach. We map each of the covariates $x^{(k)}$ and the time variable $t$ from their original scale into the interval $[0,1]$ in a way that distributes them as uniformly as possible.
Specifically, we transform the event time variable using its empirical cumulative distribution function (ECDF).
@@
where $\text{ecdf}_{t}$ is the ECDF of the event times. This transforms the event times exactly evenly spaced into $[0,1]$.
For each covariate, $x^{(k)}$, we propose a monotone increasing transformation into $[0,1]$:
$\dot{x}^{(k)}=1-\exp\left(-\frac{x^{(k)}}{c^{(k)}}\right),$ (11)
where $x^{(k)}$ denotes the value of the $k$-th covariate. The parameter $c^{(k)}$ is chosen to minimize the Kolmogorov–Smirnov (KS) test statistic for uniformity of the covariate values evaluated at the events. Specifically, by varying $c^{(k)}$, the empirical cumulative distribution function of the covariate changes and can be compared to a uniform distribution using the KS test, which evaluates the supremum of the difference between the two distributions. The KS statistic ranges from 0 to $1$, and our goal is to minimize it. Once the optimal $c^{(k)}$ is determined, the transformation in (11) is applied to both event and non-event covariate values.
Although these transformations may seem to complicate interpretation, the values can be mapped back to the original scale, allowing the results to be interpreted in terms of the original covariates, while maintaining improved numerical stability in fitting the time-varying non-linear functions.
Excluded Covariate AIC Difference LogLik Difference
Difference in prior papers -3886.01 1955.02
Author citation popularity -2858.46 1435.35
Difference in author citation popularity -408.09 215.52
Paper outdegree popularity -433.07 232.89
Prior papers -19288.83 9651.47
Prior joint papers (Linear) -3210100.35 1605046.84
Collaborate with citing author -341.42 176.77
Paper citation popularity -34990.31 17508.85
Paper pair cocitation -15622.95 7809.36
Author citation repetition -6249.01 3134.45
Cite paper and its references -19968.94 10001.37
Author self citation -33584.8 16789.89
Table 2: Contribution of individual network attributes. For each row in the table, we report the difference in AIC and log-likelihood (LogLik) between the full model (including all 12 covariates) and a reduced model that excludes the covariate listed in that row.